OpenStudy (anonymous):

find the quotient y^9 ÷y^3 y^3 y^6 y^9 y^27

jimthompson5910 (jim_thompson5910):

When dividing exponential expressions with the same base, you subtract the corresponding exponents.

jimthompson5910 (jim_thompson5910):

So what does that mean in this case?

OpenStudy (anonymous):

I'm not really sure. :/ I've been trying to figure this out for a few days

jimthompson5910 (jim_thompson5910):

Subtract the exponents: 9-3 = ???

OpenStudy (anonymous):

OH okay thank you! You actually made it make sense so it's 6 :D right?

jimthompson5910 (jim_thompson5910):

Yes, so this means \[\Large y^9 \div y^3 = y^6\]

OpenStudy (anonymous):

Thank you again :) Now I'll know how to do it next time!

jimthompson5910 (jim_thompson5910):

That's great :), you're welcome